Help me!!!
Solve the equation for y:
2x+3y=6
3x+4y=12
x-2y=8

Answer by strawberry(4)

when you either solve for x or y, just make 1 of them into 0
2x+3y=6
3y = 6
y = 2

3x+4y=12
4y = 12
y = 3

x-2y=8
-2y = 8
y = -4